# On the expressiveness and spectral bias of KANs

### v1

@@ 4.3 Gaussian random field
To approximate the Gaussian random field, we sample $f$ using the KL expansion [31]. We sample $N=5000$ points uniformly from $[-1,1]^{d}$ and calculate the (empirical) covariance matrix $K$. Then we truncate its first $m<N$ eigenpairs $\lambda_{i},\phi_{i}$, with the cutoff threshold $\lambda_{m+1}<0.1\lambda_{1}\leq\lambda_{m}$ and sample $f$ approximately via
@@
where $\xi_{i}$ are i.i.d standard Gaussians $N(0,1)$. For $f$ with different scales $\sigma$ and dimensions $d$, we split the points into $80\%$ training and $20\%$ testing points. We use MLPs and KANs with different sizes to regress on the training set, with the mean squared loss as the loss function. For MLPs, we use $500$ iterations of LBFGS iteration, and for KANs, we use the grid extension technique, with grid sizes $(10,20,30,40,50)$, each trained with $100$ iterations of LBFGS.
We plot the loss curves here and compare the losses of different scales $\sigma$ and dimensions $d$, using an MLP of $6$ layers and $256$ neurons in each hidden layer, and KANs with $10$ neurons in each hidden layer and $2,3,4$ layers; see Figure 3 for the regression loss on the training set with dimensions $2,3,4$ and scales $2^{i}$, $i=0,-1,-2,-3$. We see that for larger scale and smoother functions, MLP performs better, while for smaller scale and rough functions, KANs perform better without suffering much from spectral biases, and grid extension is especially helpful. We remark that one can choose smaller grid sizes of KANs for smoother functions and obtain more accurate regressions.
Precisely since KANs are not susceptible to spectral biases, they are likely to overfit to noises. As a consequence, we notice that KANs are more subject to overfitting on the training data regression when the task is very complicated; see the second line of Figure 4. On the other hand, we can increase the number of training points to alleviate the overfitting; see the last line of Figure 4 where we increased the number of training and test samples by $10$x. We remark that the current implementation of grid extension is prone to oscillation after refining grids during the undersampled regime, as observed in [55], and we will improve it in future works.
@@ 4.4 PDE example
In this example, we solve the 1D Poisson equation with a high-frequency solution, similar to [75]. To be precise, consider the equation with zero Dirichlet boundary condition
@@
The different frequencies are normalized in a way that for $k>1$, the ground truth has the same energy ($H^{1}$) norm. We use the variational form of the elliptic equation and the associated Deep Ritz Method [83]. Parametrizing $u$ by a neural network, we minimize the loss
@@
For frequencies $k=2,4,8,16,32$, we use $2000$ uniformly spaced sample points and the neural network using an MLP of 6 layers and 256 neurons in each hidden layer and a KAN of 2 layers with 10 neurons in the hidden layer. We choose the hyperparameter $\lambda=0.01$ balancing the energy and boundary loss and perform LBFGS iterations. For MLPs, we use 200 iterations, and for KANs, we use grid sizes $(20,40)$, each trained with 100 iterations. We plot the relative $L^{2}$ and $H^{1}$ losses compared to the ground truth in Figure 5. We can see that KANs perform consistently better, and the residue barely deteriorates when the frequency increases, whereas it becomes extremely hard for MLPs to optimize when $k=16,32.$
@@ 5 Concluding Remarks
Appendix A Proofs
@@
The proof follows from the fact that each KAN activation function $\phi_{l,i,j}$ can be represented by a ReLUk neural network with one hidden layer of width $G+2k+1$. This is due to the fact that each $\phi_{l,i,j}$ is a linear combination of $B$-splines (since we are assuming that $w_{b}=0$ in (4)). This implies that it must be a compactly support piecewise polynomial spline of degree $k$ whose $(k-1)$-st derivative changes only on the extended grid which has $G+2k+1$ points. We can now easily construct a shallow ReLUk network with one hidden node per grid point which matches the function $\phi_{l,i,j}$. ∎
@@
It is well-known that the Gram matrix $C$ is well-conditioned uniformly in $G$ for a fixed $k$, i.e. $\lambda_{G+k-1}(C)/\lambda_{1}(C)\leq K$ for a fixed constant $K$ depending only upon $k$. See for instance [22], Theorem 4.2 in Chapter 5, where it is shown that the $L_{2}$-norm of a spline and the properly scaled $\ell_{2}$-norm of its B-spline coefficients are equivalent up to a constant depending only on $k$. This is equivalent to the well-conditioning of the Gram matrix $C$.

### v2

@@ 4.3 Gaussian random field
To approximate the Gaussian random field, we sample $f$ using the KL expansion Karhunen (1947). We sample $N=5000$ points uniformly from $[-1,1]^{d}$ and calculate the (empirical) covariance matrix $K$. Then we truncate its first $m<N$ eigenpairs $\lambda_{i},\phi_{i}$, with the cutoff threshold $\lambda_{m+1}<0.1\lambda_{1}\leq\lambda_{m}$ and sample $f$ approximately via
@@
where $\xi_{i}$ are i.i.d standard Gaussians $N(0,1)$. For $f$ with different scales $\sigma$ and dimensions $d$, we split the points into $80\%$ training and $20\%$ testing points. We use MLPs and KANs with different sizes to regress on the training set, with the mean squared loss as the loss function. For MLPs, we use $500$ iterations of LBFGS iteration, and for KANs, we use the grid extension technique, with grid sizes $(10,20,30,40,50)$, each trained with $100$ iterations of LBFGS. We can see that ReLUk MLPs performs even worse than a standard MLP.
We plot the loss curves here and compare the losses of different scales $\sigma$ and dimensions $d$, using an MLP of $256$ neurons in each hidden layer and optimized over its depth, and KANs with $10$ neurons in each hidden layer and $2,3,4$ layers; see Figure 2 for the regression loss on the training set with dimensions $2,3,4$ and scales $2^{i}$, $i=0,-1,-2,-3$. We see that for larger scale and smoother functions, MLP performs better, while for smaller scale and rough functions, KANs perform better without suffering much from spectral biases, and grid extension is especially helpful. We remark that one can choose bigger grid sizes of KANs for smoother functions and obtain more accurate regressions.
Precisely since KANs are not susceptible to spectral biases, they are likely to overfit to noises. As a consequence, we notice that KANs are more subject to overfitting on the training data regression when the task is very complicated; see the second line of Figure 3. On the other hand, we can increase the number of training points to alleviate the overfitting; see the last line of Figure 3 where we increased the number of training and test samples by $10$x. We remark that the current implementation of grid extension is prone to oscillation after refining grids during the undersampled regime, as observed in Rigas et al. (2024), and we will improve it in future works.
@@ 4.4 PDE example
In this example, we solve the 1D Poisson equation with a high-frequency solution, similar to Xu et al. (2019a). To be precise, consider the equation with zero Dirichlet boundary condition
@@
The different frequencies are normalized in a way that for $k>1$, the ground truth has the same energy ($H^{1}$) norm. We use the variational form of the elliptic equation and the associated Deep Ritz Method Yu et al. (2018). Parametrizing $u$ by a neural network, we minimize the loss
@@
For frequencies $k=2,4,8,16,32$, we use $2000$ uniformly spaced sample points and the neural network using an MLP of 6 layers and 256 neurons in each hidden layer and a KAN of 2 layers with 10 neurons in the hidden layer. We choose the hyperparameter $\lambda=0.01$ balancing the energy and boundary loss and perform LBFGS iterations. For MLPs, we use 200 iterations, and for KANs, we use grid sizes $(20,40)$, each trained with 100 iterations. We plot the relative $L^{2}$ and $H^{1}$ losses compared to the ground truth in Figure 4. We can see that KANs perform consistently better, and the residue barely deteriorates when the frequency increases, whereas it becomes extremely hard for MLPs to optimize when $k=16,32.$ We refer a 2D case to the appendix A.
@@ 5 Concluding Remarks
Appendix A 2D Poisson equation with high frequency components
In this experiment, we solve a 2D Poisson equation with a high-frequency solution using the Deep Ritz method. The equation with zero Dirichlet boundary condition has the form
$-u_{xx}-u_{yy}=f\,\,\,\,\text{in}[0,1]^{2}\,,\quad u(x,0)=u(x,1)=u(0,y)=u(1,y)=0\,.$ (14)
with a right-hand side and a corresponding exact solution depending on the frequency $k$ as
$f=2\pi^{2}\sin(\pi x)\sin(\pi y)+2\pi^{2}{k}\sin(k\pi x)\sin(k\pi y),\quad u=\sin(\pi x)\sin(\pi y)+\frac{1}{k}\sin(k\pi x)\sin(k\pi y).$
For frequencies $k=2,4,16$, we use a uniform grid with $101$ sample points per direction with a MLP of 6 layers and 256 neurons in each hidden layer and a KAN of 2 layers with 10 neurons in the hidden layer. We use LBFGS for 50 iterations and plot the loss compared to the ground truth in Figure 5. We observe similarly as in the 1D case that KANs outform MLPs when solving equation with multiple scales.
More importantly, MLPs cannot learn the correct shape, even with $k=4$ after 100 iterations, whereas KANs can learn a solution quite close to the ground truth with just $5$ iterations for $k=4$ and $10$ iterations for $k=16$; see Figure 6.
Figure 5: 2D Poisson. Losses of MLP and KAN with different frequencies of the solution.
Figure 6: 2D Poisson. True solution and neural net solutions.
Appendix B 1D waves of increasing amplitudes
Appendix C Proofs
@@
The proof follows from the fact that each KAN activation function $\phi_{l,i,j}$ can be represented by a ReLUk neural network with one hidden layer of width $G+2k+1$. This is due to the fact that each $\phi_{l,i,j}$ is a linear combination of $B$-splines (since we are assuming that $w_{b}=0$ in (4)). This implies that it must be a compactly support piecewise polynomial spline of degree $k$ whose $(k-1)$-st derivative changes only on the extended grid which has $G+2k+1$ points. Based upon this, we can construct a shallow ReLUk network with one hidden node per grid point which matches each function $\phi_{l,i,j}$ (see for instance Xu (2020)). Since there are a total of $W^{2}$ such functions in each layer, this means that we can implement a single KAN layer using a shallow ReLUk network of width $(G+2k+1)W^{2}$. By composing these networks, we obtain the desired MLP representing the KAN. We remark that after grid extension, the grid points for each for each of the spline functions $\phi_{l,i,j}$ may become different. For this reason, it is not clear whether the width in this construction can be reduced. ∎
@@
It is well-known that the Gram matrix $C$ is well-conditioned uniformly in $G$ for a fixed $k$, i.e. $\lambda_{G+k-1}(C)/\lambda_{1}(C)\leq K$ for a fixed constant $K$ depending only upon $k$. See for instance DeVore & Lorentz (1993), Theorem 4.2 in Chapter 5, where it is shown that the $L_{2}$-norm of a spline and the properly scaled $\ell_{2}$-norm of its B-spline coefficients are equivalent up to a constant depending only on $k$. This is equivalent to the well-conditioning of the Gram matrix $C$.